OpenStudy (anonymous):

Given that x has a Poisson Distribution with u(mean)=9, what is the probability that x = 3

OpenStudy (anonymous):

P(x) = (mean)^x exp(-mean) / x! P(3) = (9)^3 exp(-9)/ 3!

OpenStudy (anonymous):

I don't understand that at all

OpenStudy (anonymous):

Neat thing about the Poisson is that from P(0) = exp(-mean) you can make inferences about the mean from observing no events during the period studied.

OpenStudy (anonymous):

so I still don't know how to get this answer

OpenStudy (tkhunny):

You get the answer by learning the structure of the Poisson Distribution and calculating it directly. This is what DWC is showing you.

OpenStudy (tkhunny):

The structure: \(P(x) = \dfrac{\lambda^{x}\cdot e^{-\lambda}}{x!}\) The calculation: \(\lambda = 9\) \(x = 3\) \(P(3) = \dfrac{9^{3}\cdot e^{-9}}{3!}\) Do NOT say you don't know how to do that. If you don't know how to do that, then you do not have the right prerequisites to be in this class. Let's do better.